# thanks: Full author list given at the end of the paper

### v2

@@ III.3 Fit of the joint angular distributions
where $\vec{\xi}_{i}$ denotes the kinematic observables, including the helicity angles and Daltiz-plot variables, while $\vec{\eta}$ denotes the free parameters to be determined. The $M(\vec{\xi}_{i};\vec{\eta})$ is the total decay amplitude, and $\epsilon(\vec{\xi}_{i})$ is the detection efficiency, parameterized in terms of the kinematic variables $\vec{\xi}_{i}$. It is important to note that all $\vec{\xi}_{i}$ are calculated using four-momentum updated by kinematic fit, which constrains the invariant mass of all resonances in the signal decay to their world average masses [41]. The background contribution to the joint likelihood is subtracted based on the calculated likelihoods using data in $M_{\text{BC}}\in[2.250,2.270]~\text{GeV}/c^{2}$. For the mis-reconstructed signal events, the contribution to the likelihood is based on the signal MC simulation. The integration of the normalization factor is calculated with a large PHSP signal simulation samples as $\int\epsilon(\vec{\xi}_{i})|M(\vec{\xi}_{i};\vec{\eta})|^{2}\text{d}\vec{\xi}_{i}=\frac{1}{N_{\text{gen}}}\sum_{k_{\text{sim}}}^{N_{\text{sim}}}|M(\vec{\xi}_{k_{\text{sim}}};\vec{\eta})|^{2}$, where $N_{\text{gen}}$ is the total number of the generated events, $N_{\text{sim}}$ is the number of the events surviving all selection criteria and $k_{\text{sim}}$ is the event index.
@@
The negative log-likelihood with background subtraction over the five signal channels and data samples collected from thirteen energy points is minimized using the Minuit [74] algorithms. The decay asymmetry parameters of the $\Lambda$, $\Sigma^{+}$, and $\Sigma^{0}$ baryons are fixed according to the PDG [41]. In the nominal fit, the parameters $\alpha_{BP}$ and $\Delta_{BP}$ are set to $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$, respectively, in accordance with the assumption of CPC\!P conservation. Here, $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$ represent the parameters associated with the antiparticle $\bar{\Lambda}_{c}^{-}$ decays. The results of the fit are presented in Fig. 4. An alternative test is also performed where the assumption of CPC\!P conservation are relaxed and $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$ are treated as free parameters in the fit.
@@
Figure 4: Fit results of helicity angles combining all energy points. They are angular distribution about $pK_{S}^{0}$, $pK^{-}\pi^{+}$, $\Lambda\pi^{+}$, $\Sigma^{+}\pi^{0}$, and $\Sigma^{0}\pi^{+}$, respectively. Black points with error bars are data; the yellow shaded region indicates combinatorial background events and black shaded region is the mis-reconstructed signal events.
@@ III.4 Systematic uncertainties
  • Selection criteria. The systematic uncertainties regarding selection criteria primarily arise from the reconstruction of final states, $\Delta E$ requirement, and $M_{\text{BC}}$ signal selection. The reconstruction efficiency of proton, pion, and kaon tracks is studied with control samples of $^{+}e^{-}\to p\bar{p}\pi^{+}\pi^{-}, e+e$$^{+}e^{-}\to\pi^{+}\pi^{-}\pi^{+}\pi^{-}$, and $J/\psi\to K^{0}_{S}K^{\pm}\pi^{\mp}$ decays, respectively. The $\pi^{0}$ meson reconstruction efficiency is studied with the $\psi(3686)\to\pi^{0}\pi^{0}J/\psi$ and $e^{+}e^{-}\to\omega\pi^{0}$ decays. The control channels for the $K^{0}_{S}$ reconstruction efficiency are the $J/\psi\to\phi K^{0}_{S}K^{\mp}\pi^{\pm}$ and $J/\psi\to K^{*}(892)^{\pm}K^{\mp}$ decays. For the reconstruction efficiency of the $\Lambda$ baryon the $J/\psi\to\bar{p}K^{+}\Lambda$ and $J/\psi\to\Lambda\bar{\Lambda}$ decay samples are used [75]. The difference in reconstruction efficiency for photons between data and MC simulation is studied with the $J/\psi\to\pi^{+}\pi^{-}\pi^{0}$ decay and found to be $0.5\%$ and $1.5\%$ in the barrel and endcap of the EMC. A randomly generated efficiency correction factor is used to adjust the PHSP simulation. Requirements relating to the MUC depth, probability of electron, and pion assumption are studied with a control sample of $\Lambda^{+}_{c}\to\Lambda\pi^{+}$ decays. The potential differences in event selection efficiencies for data and MC simulation are taken into account in the fit described in the previous section via reweighting the PHSP sample. The uncertainties related to the $\Delta E$ and $M_{\text{BC}}$ requirements are evaluated by smearing the signal simulation with a Gaussian resolution function. The changes in the fit results based on new survived signal simulation are taken to be systematic uncertainties.
  • Performance of truth matching. The maximum difference between the matching condition with 9 or 11 degree and the nominal result is considered as systematic uncertainty.
@@
  • Signal components. The amplitude-model related uncertainties are computed using the alternative models in Ref. [73]. With alternative polarimeter field inputs, the angular fit is repeated and the standard deviations and difference of mean value are assigned to the model uncertainties. In Ref. [72], the statistical uncertainties are estimated by varying the parameter values in the default model. These parameters are sampled from a Gaussian function, whose $\mu$ and $\sigma$ are given by the central and error values from the default model fit results in Ref. [73]. The polarimeter vector fields are reproduced for each varied parameter set. The angular fit procedure is repeated using the corresponding inputs, and the statistical component is given by taking the standard deviations and the difference of the mean value.
@@
  • Background components. The background shape is extracted from the data events in the $M_{\text{BC}}$ sideband regions. To estimate the possible difference between the $M_{\text{BC}}$ sideband and signal regions, a weight factor is derived by comparing the $M(K^{-}\pi^{+})^{2}$ distributions in the $M_{\text{BC}}$ sideband and signal regions of inclusive simulation. Data events in the $M_{\text{BC}}$ sideband region are reweighted and the fit procedure is repeated. The difference is then taken as the uncertainty from background model. Background fractions are fixed according to the fit results from the nominal fit. To estimate systematic uncertainties, the background fraction for each energy point is varied by $\pm 1\sigma$ and the fit procedure is repeated. The quadratic sum of the largest variation from each fraction are assigned as systematic uncertainties from the background fraction. The mis-reconstructed model is examined using an alternative signal simulation produced with new input parameters $\alpha_{\Sigma^{0}\pi^{+}}$ and $\alpha_{\Sigma^{+}\pi^{0}}$ which are changed within $\pm 1\sigma$ of the results presented here. The uncertainty due to the combinational background model is estimated by varying the relative weights between $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ pairs and other hadronic events based on the uncertainties of their cross-section ratio. The relevant systematic uncertainties from background size are examined by repeating the fits with an alternative background size obtained from the Gaussian sampling of the fitted parameters. The series of fitted parameters, assumed to follow a Gaussian distribution, are used to estimate the systematic uncertainty for the background size. The uncertainty is determined by summing the fitted Gaussian resolution and the deviation of the fitted Gaussian mean from the nominal value.
@@
  • Fixed parameters. The systematic uncertainty due to the input parameters is evaluated by varying these parameters using a Gaussian sampling method. For each parameter, the obtained results are expected to follow a Gaussian distribution. The sum of the fitted Gaussian resolution and the difference between the fitted Gaussian mean and the nominal result is taken as the systematic uncertainty.
@@ IV.1 Transverse polarization
The results of the angular distribution fit, the $\alpha_{0}$, the phase difference $\Delta\Phi$, its significance, as well as the largest transverse polarization $\mathcal{P}_{y}^{max}$, are summarized in Table 3. The EMFF-related $\alpha_{0}$ is consistent with previous results but more precise. The ratio of $|G_{E}/G_{M}|$ is extracted using the equation $|G_{E}/G_{M}|^{2}(1-v^{2})=(1+\alpha_{0})/(1-\alpha_{0})$, where $v$ is the velocity of the $\Lambda^{+}_{c}$ in CM system, normalized to the speed of light. The ratios are compared with the previous BESIII measurements and theoretical predictions, as shown in Fig. 5. The measurements of $\sin\Delta\Phi$ are also presented in Fig. 5, with the theoretical predictions [76, 77] and other experimental results [15, 78]. There is some inconsistency between the measurements and the theoretical predictions.
Table 3: Results for $\alpha_{0}$, $\Delta\Phi$, and transverse polarization of $\Lambda_{c}^{+}$ at 13 energy points, where the first uncertainties are statistical and the second are systematic. In the table, SLSL means the significance level of non-zero $\Delta\Phi$, and $\mathcal{P}_{y}^{max}$ is the largest transverse polarization.
$\sqrt{s}~[\text{MeV}]$ $\alpha_{0}$~in previous work [15, 78] $\alpha_{0}$~in this work $\Delta\Phi~[\text{rad}]$ SLSL $\mathcal{P}_{y}^{max}$
@@
Figure 5: Comparison of $|G_{E}/G_{M}|$ and $\sin\Delta\Phi$ between prediction and this work. Red points with error bars correspond to this work, and the gray ones are previous measurement results from BESIII [15, 78]. The purple dashed line with sky blue shadow is predicted from Ref. [76], and the blue solid line represents the predicted results with Ref. [77].
Of particular interest is the fact that the $\Lambda_{c}^{+}$ baryon is significantly polarized with respect to the normal of the production plane. The size of polarization is strongly dependent on $\Lambda_{c}^{+}$ momentum direction $\cos\theta_{0}$ with a maximum value of approximately 12%. To illustrate the polarization effect, the moment $\textless\sin\theta_{1}\sin\phi_{1}\textgreater$ is calculated as the average of $\sin\theta_{1}\sin\phi_{1}$ within a certain range of $\theta_{0}$. It is expected to be proportional to $\alpha_{BP}P_{y}(\cos\theta_{0})/(1+\alpha_{0}\cos^{2}\theta_{0})$. We divide the range of $\cos\theta_{0}$, (1,1)(-1,1), into eight bins, and calculate the moment for both data and reweighted simulation at energy points where the significance of $\Delta\Phi$ is larger than $5\sigma$, as shown in Fig. 6. To compute the significance, we perform a floating fit for $\Delta\Phi$ and a fixed fit for $\Delta\Phi=0$ to obtain the likelihood values and the degrees of freedom, from which we derive the significance of non-zero $\Delta\Phi$. The sign($\alpha_{BP}$) is used to avoid cancellation between charge-conjugated processes. The significance is reported in Table 3.
Figure 6: The moment $\textless\text{sign}(\alpha_{BP})\sin\theta_{1}\sin\phi_{1}\textgreater$ distribution to show the effect of tranverse polarization. It depends on $\cos\theta_{0}$ for $\Lambda_{c}^{+}\to pK^{0}_{S}$, $\Lambda_{c}^{+}\to\Lambda\pi^{+}$, and $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$ channels and is drawn using points with significance of $\Delta\Phi$ larger than $5\sigma$. Points with error bars correspond to data and red band represents the simulation reweighted according to the fit results with total uncertainties.
@@ IV.2 Decay asymmetry and SM test
As the SM predicts no CPC\!P asymmetry in these decay channels, we assume CPC\!P symmetry, meaning the decay asymmetry parameters $\alpha_{BP}=-\bar{\alpha}_{BP}$ and $\Delta_{BP}=-\bar{\Delta}_{BP}$ in the angular distribution analysis. In addition to the transverse polarization, the decay asymmetry parameters are obtained and shown in Fig. 7. In the SM, the amplitude for a spin-$1/2$ baryon decaying into a spin-$1/2$ baryon and a spin-0 meson can be written as $\mathcal{M}=i\bar{u}_{f}(A-B\gamma_{5})u_{i}$, where $A$ and $B$ are constants, that reflect the $S$ and $P$-wave amplitudes with different orbital angular momentum in weak decays, and directly reflect the mechanism of different weak decays. The $u_{i}$ and $\bar{u}_{f}$ are spinors describing the initial and final baryons [19]. The calculated method of $|A|$ and $|B|$ has been applied in $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ [28], successfully promoting the development of effective models [30, 31, 33]. In this work, $A$ and $B$ can also be derived, and their two-dimensional displays are shown in Fig. 8. We provide only the density plots and do not present numerical results due to the nonlinear issues encountered during the calculation, which makes it difficult to provide reliable uncertainty estimates.
@@
Figure 7: The comparison about predictions and measurements of decay asymmetry parameters. Gray points are theoretical predictions [17, 19, 20, 21, 23, 22, 24, 18, 25, 34, 35, 36, 37, 29, 30, 31, 33] and experimental measurements [16, 38, 39, 40, 15, 26, 27, 32, 41] are marked in blue. Red points and yellow shadows represent the results obtained in this work. In the legend, the names of the theoretical studies and experimental measurements match the colors used in the figures, arranged in chronological order. For each theoretical study, the main theoretical models used are also specified.
@@
Figure 8: The two-dimensional density of $|A|$ and $|B|$. Markers are theoretical predictions [19, 20, 21, 18, 35, 37, 34, 36, 31, 33] and experimental measurements [39], and shadows representing the distribution of $|A|$ and $|B|$ obtained in this work, $G_{F}$ is Fermi coupling constant, $c$ and $d$ means whether the model uses the results from Belle experiment with the BFs and decay asymmetry parameters of $\Xi_{c}^{0}\to\Xi^{0}\pi^{0}$, $\Xi^{0}\eta$, and $\Xi^{0}\eta^{\prime}$. The contours correspond to 68.2%, 95.4%, 99.7% confidence levels, respectively.
@@
A secondary fit is also performed where the CPC\!P constraint is removed and is used to test the SM. The $\alpha$-induced CPC\!P observable, weak and strong phase shifts are constructed in $A^{\alpha_{BP}}_{C\!P}=\frac{\alpha_{BP}+\bar{\alpha}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, $\text{tan}\phi_{C\!P}=\frac{\beta_{BP}+\bar{\beta}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, $\text{tan}\Delta_{S}=\frac{\beta_{BP}-\bar{\beta}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, can be extracted for $\Lambda_{c}^{+}$ decay channels, where $\beta_{BP}=\sqrt{1-\alpha_{BP}^{2}}\sin\Delta_{BP}$, $\phi_{C\!P}$ and $\Delta_{S}$ denote the weak interaction phase shift between $P$-wave and $S$-wave introduced by phase angle in the Cabibbo-Kobayashi-Maskawa (CKM) matrix and the strong phase shift. The results are summarized in Table 4.
@@
Table 4: Results of decay observables, where $\langle\alpha_{BP}\rangle$, $\langle\Delta_{BP}\rangle$, and $\delta_{p}-\delta_{s}$ are measured based on the assumption of CPC\!P conservation, while others are under the assumption of CPC\!P violation. Some negligible terms are denoted by “”. All angular values in the table are expressed in radians.
Parameter $\Lambda_{c}^{+}\to pK_{S}^{0}$ $\Lambda_{c}^{+}\to\Lambda\pi^{+}$ $\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ $\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$

### v3

@@ III.3 Fit of the joint angular distributions
where $\vec{\xi}_{i}$ denotes the kinematic observables, including the helicity angles and Dalitz-plot variables, while $\vec{\eta}$ denotes the free parameters to be determined. The $M(\vec{\xi}_{i};\vec{\eta})$ is the total decay amplitude, and $\epsilon(\vec{\xi}_{i})$ is the detection efficiency, parameterized in terms of the kinematic variables $\vec{\xi}_{i}$. It is important to note that all $\vec{\xi}_{i}$ are calculated using four-momenta updated by a kinematic fit, which constrains the invariant mass of all resonances in the signal decay to their world average masses [ParticleDataGroup:2024cfk]. The integration of the normalization factor is calculated with a large PHSP signal simulation samples as $\int\epsilon(\vec{\xi}_{i})|M(\vec{\xi}_{i};\vec{\eta})|^{2}\text{d}\vec{\xi}_{i}=\frac{1}{N_{\text{gen}}}\sum_{k_{\text{surv}}}^{N_{\text{surv}}}|M(\vec{\xi}_{k_{\text{surv}}};\vec{\eta})|^{2}$, where $N_{\text{gen}}$ is the total number of the generated events, $N_{\text{surv}}$ is the number of the events surviving all selection criteria and $k_{\text{surv}}$ is the event index. The background contribution to the joint likelihood is subtracted based on the calculated likelihoods using data in $M_{\text{BC}}\in[2.250,2.270]~\text{GeV}/c^{2}$. For the mis-reconstructed signal events, the contribution to the likelihood is based on the signal MC simulation. Both components are scaled to match the corresponding yields in data, and the scale factors can be obtained from the $M_{\text{BC}}$ fitting. The likelihood function with background subtraction can be expressed as
$\mathcal{L}_{\text{sig}}=K\times(\mathcal{L}_{\text{data}}-\mathcal{L}_{\text{bkg}}-\mathcal{L}_{\text{mis}}),$ (7)
where the factor $K$ is used to correct uncertainties in unbinned maximum likelihood fit as suggested in Refs. [BESIII:2022udq, Langenbruch:2019nwe]. The subscripts “bkg” and “mis” correspond to the background and mis-reconstructed components, respectively.
@@
The negative log-likelihood with background subtraction over the five signal channels and data samples collected from thirteen energy points is minimized using the Minuit [James:1975dr] algorithms. The decay asymmetry parameters of the weak decays $\Lambda\to p\pi^{-}$ and $\Sigma^{+}\to p\pi^{0}$ and their charge conjugated processes $\bar{\Lambda}\to\bar{p}\pi^{+}$ and $\bar{\Sigma}^{-}\to\bar{p}\pi^{0}$ are fixed according to the experimentally measured decay asymmetry parameters [BESIII:2022qax, ParticleDataGroup:2024cfk]. In the nominal fit, the parameters $\alpha_{BP}$ and $\Delta_{BP}$ are set to $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$, respectively, in accordance with the assumption of CPC\!P conservation. Here, $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$ represent the parameters associated with the antiparticle $\bar{\Lambda}_{c}^{-}$ decays. The results of the fit are presented in Fig. 5, and the full correlation matrix is provided in Table A1 of Appendix A. An alternative test is also performed where the assumption of CPC\!P conservation are relaxed and $\bar{\alpha}_{BP}$ and $\bar{\Delta}_{BP}$ are treated as free parameters in the fit.
@@
Figure 5: Fit results of helicity angles combining all energy points. They are angular distribution about $pK_{S}^{0}$, $pK^{-}\pi^{+}$, $\Lambda\pi^{+}$, $\Sigma^{+}\pi^{0}$, and $\Sigma^{0}\pi^{+}$, respectively. Black points with error bars are data; the green shaded region indicates combinatorial background events and black shaded region is the mis-reconstructed signal events.
@@ III.4 Systematic uncertainties
[…]
  • Performance of truth matching. In the nominal measurement, the angle between the reconstructed and true photon momentum directions from simulations are required to be below 10 degrees. To evaluate the effect due to this truth matching requirement, two different matching conditions of 9 and 11 degrees are tested and the resultant largest difference from the nominal result is considered as systematic uncertainty.
@@
  • Signal components. The amplitude-model related uncertainties are computed using the alternative models in Ref. [LHCb:2022sck]. With alternative $\Lambda^{+}_{c}$ polarimetry, the angular fit is repeated and the standard deviations and difference of mean value are assigned to the model uncertainties. In Ref. [LHCb:2023crj], the statistical uncertainties are estimated by varying the parameter values in the default model. These parameters are sampled from a Gaussian function, whose $\mu$ and $\sigma$ are given by the central and error values from the default model fit results in Ref. [LHCb:2022sck]. The polarimeter vector fields are reproduced for each varied parameter set. The angular fit procedure is repeated using the corresponding inputs, and the statistical component is given by taking the standard deviations and the difference of the mean value.
@@
  • Background components. Background fractions are fixed according to the fit results from the nominal fit. To estimate systematic uncertainties, the background fraction for each energy point is varied by $\pm 1\sigma$ and the fit procedure is repeated. The quadratic sum of the largest variation from each fraction are assigned as systematic uncertainties from the background fraction. The mis-reconstructed model is examined using an alternative signal simulation produced with new input parameters $\alpha_{\Sigma^{0}\pi^{+}}$ and $\alpha_{\Sigma^{+}\pi^{0}}$ which are changed within $\pm 1\sigma$ of the results presented here. The background shape is extracted from the data events in the $M_{\text{BC}}$ sideband regions. For $\Lambda_{c}^{+}\to pK^{-}\pi^{+}$ channel, to estimate the possible difference between the $M_{\text{BC}}$ sideband and signal regions, a weight factor is derived by comparing the $M(K^{-}\pi^{+})^{2}$ distributions in the $M_{\text{BC}}$ sideband and signal regions of inclusive simulation, since the most significant discrepancy appears in the $M(K^{-}\pi^{+})^{2}$ distribution, while the others show consistent behavior. Data events in the $M_{\text{BC}}$ sideband region are corrected by the weight factor and the fit procedure is repeated. The difference is then taken as the uncertainty from background model. For two-body decays, the uncertainty due to the combinational background model is estimated by varying the relative weights between $\Lambda_{c}^{+}\bar{\Lambda}_{c}^{-}$ pairs and other hadronic events based on the uncertainties of their cross-section ratio. The relevant systematic uncertainties from background size are examined by repeating the fits with an alternative background size obtained from the Gaussian sampling of the fitted parameters. The series of fitted parameters, assumed to follow a Gaussian distribution, are used to estimate the systematic uncertainty for the background size. The fitted Gaussian means are found to be consistent to the nominal result, typically with central value deviating at the level of $10^{-3}$. To be conservative, the corresponding systematic uncertainty is determined by summing the fitted Gaussian width and the deviation of the fitted Gaussian mean from the nominal value.
@@
  • Fixed parameters. The systematic uncertainty due to the input parameters is evaluated by varying these parameters using a Gaussian sampling method. For each parameter, the obtained results are expected to follow Gaussian distributions. The sum of the fitted Gaussian resolution and the difference between the fitted Gaussian mean and the nominal result is taken as the systematic uncertainty. The fitted means are close to zero [Zhong:2024qqs], but they are included with the width as a conservative estimation.
@@ IV.1 Transverse polarization
The results of the angular distribution fit, the $\alpha_{0}$, the phase difference $\Delta\Phi$, its significance, as well as the largest transverse polarization $\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y,max}}$, are summarized in Table 3. The EMFF-related $\alpha_{0}$ is consistent with previous results but more precise. The ratios $|G_{E}/G_{M}|$ are compared with the previous BESIII measurements [Chen:2023oqs] and theoretical calculations [Wan:2021ncg, Chen:2023oqs], as shown in Fig. 6. The measurements of $\sin\Delta\Phi$ are also presented with the theoretical calculations [Wan:2021ncg, Chen:2023oqs], which significantly differs from the theoretical predictions, providing information beyond current theoretical expectations.
Table 3: Results for $\alpha_{0}$, $\Delta\Phi$, and transverse polarization of $\Lambda_{c}^{+}$ at 13 energy points, where the first uncertainties are statistical and the second are systematic. In the table, SLSL means the significance level of non-zero $\Delta\Phi$, $\sigma$ represents the standard deviation, and $\bm{{\mathcal{P}_{\Lambda_{c}^{+}}^{y,max}}}$ is the largest transverse polarization.
$\sqrt{s}~[\text{MeV}]$ $\alpha_{0}$~in previous work [BESIII:2019odb, BESIII:2023rwv] $\alpha_{0}$~in this work $\Delta\Phi~[\text{rad}]$ SLSL $\bm{\mathcal{P}_{\Lambda_{c}^{+}}^{y,max}}$
@@
Figure 6: Comparison of $|G_{E}/G_{M}|$ and $\sin\Delta\Phi$ between theoretical calculation and this work. Red points with error bars correspond to this work, and the gray ones are previous measurement results from BESIII [BESIII:2019odb, BESIII:2023rwv]. The purple dashed line with sky blue shadow is calculated from Ref. [Wan:2021ncg], and the blue solid line represents the calculated results with Ref. [Chen:2023oqs].
Of particular interest is the fact that the $\Lambda_{c}^{+}$ baryon is significantly polarized with respect to the normal of the production plane. The size of polarization is strongly dependent on $\Lambda_{c}^{+}$ momentum direction $\cos\theta_{0}$ with a maximum value of approximately 12%. To illustrate the polarization effect, the moment $\textless\sin\theta_{1}\sin\phi_{1}\textgreater$ is calculated as the average of $\sin\theta_{1}\sin\phi_{1}$ within a certain range of $\cos\theta_{0}$. It is expected to be proportional to $\alpha_{BP}P_{y}(\cos\theta_{0})/(1+\alpha_{0}\cos^{2}\theta_{0})$. We divide the range of $\cos\theta_{0}$, (1,1)(-1,1), into eight bins, and calculate the moment for experimental data and provide the best fit curve at all energy points, as shown in Fig. 7. To compute the significance, we perform a floating fit for $\Delta\Phi$ and a fixed fit for $\Delta\Phi=0$ to obtain the likelihood values and the degrees of freedom, from which we derive the significance of non-zero $\Delta\Phi$. The sign($\alpha_{BP}$) is used to avoid cancellation between charge-conjugated processes. The significance is reported in Table 3.
Figure 7: The moment $\textless\text{sign}(\alpha_{BP})\sin\theta_{1}\sin\phi_{1}\textgreater$ distribution to show the effect of transverse polarization. Points with error bars correspond to data with background subtracted and the red solid line denote the global expected dependence on $\cos\theta_{0}$. The blue dashed line represents the distribution without polarization uniformly distributed in phase space.
@@ IV.2 Decay asymmetry and SM test
As the SM predicts no CPC\!P asymmetry in these decay channels, we assume CPC\!P symmetry, meaning the decay asymmetry parameters $\alpha_{BP}=-\bar{\alpha}_{BP}$ and $\Delta_{BP}=-\bar{\Delta}_{BP}$ in the angular distribution analysis. In addition to the transverse polarization, the decay asymmetry parameters are obtained and shown in Fig. 8. In the SM, the amplitude for a spin-$1/2$ baryon decaying into a spin-$1/2$ baryon and a spin-0 meson can be written as $\mathcal{M}=i\bar{u}_{f}(A-B\gamma_{5})u_{i}$, where $A$ and $B$ are constants, that reflect the $S$ and $P$-wave amplitudes with different orbital angular momentum in weak decays, arising from the parity-violating and parity-conserving components of the weak interaction, respectively. The $u_{i}$ and $\bar{u}_{f}$ are spinors describing the initial and final baryons [Xu:1992vc]. By combining the measured parameter $\alpha_{BP}$, $\Delta_{BP}$, and the branching fraction $\mathcal{B}$, a set of equations can be established to relate the observables to the decay amplitudes. These equations allow the extraction of the magnitudes of $|A|$ and $|B|$ as well as the relative phase shift $\delta_{p}-\delta_{s}$:
[…]
where $\vec{p}_{c}$ is the momentum of $B$ in the rest frame of $\Lambda_{c}^{+}$ and the auxiliary parameter $\kappa$ is defined as $\kappa=|\vec{p}_{c}|/(E_{B}+m_{B})=\sqrt{(E_{B}-m_{B})/(E_{B}+m_{B})}$. The symbol $\tau_{\Lambda_{c}^{+}}$ denotes the mean lifetime of the baryon, and $E$, $m$ represent the energy in the $\Lambda_{c}^{+}$ rest frame and the invariant mass, respectively. The calculated method of $|A|$ and $|B|$ has been applied in $\Lambda_{c}^{+}\to\Xi^{0}K^{+}$ [BESIII:2023wrw], successfully promoting the development of effective models [Zhong:2024zme, Zhong:2024qqs, Cheng:2024lsn]. In this work, $|A|$ and $|B|$ can also be derived, and their two-dimensional displays are shown in Fig. 9. We provide only the density plots and do not present numerical results due to the complex shapes of the contours cannot be well represented by a simple confidence interval.
@@
Figure 8: The comparison about theoretical calculations and experimental measurements of decay asymmetry parameters. Gray points are theoretical calculations [Korner:1992wi, Xu:1992vc, Cheng:1991sn, Cheng:1993gf, Zenczykowski:1993hw, Zenczykowski:1993jm, Datta:1995mn, Ivanov:1997ra, Sharma:1998rd, Geng:2019xbo, Zou:2019kzq, Zhong:2022exp, Liu:2023dvg, Geng:2023pkr, Zhong:2024zme, Zhong:2024qqs, Cheng:2024lsn] and experimental measurements [CLEO:1990unw, ARGUS:1991yzs, CLEO:1995qyd, FOCUS:2005vxq, BESIII:2019odb, Belle:2022uod, Belle:2022bsi, LHCb:2024tnq, ParticleDataGroup:2024cfk] are marked in blue. Red points and yellow shadows represent the results obtained in this work. In the legend, the names of the theoretical studies and experimental measurements match the colors used in the figures, arranged in chronological order. For each theoretical study, the main theoretical models used are also specified.
@@
Figure 9: The two-dimensional density of $|A|$ and $|B|$. Markers are theoretical calculations [Xu:1992vc, Cheng:1991sn, Cheng:1993gf, Ivanov:1997ra, Zou:2019kzq, Liu:2023dvg, Geng:2019xbo, Zhong:2022exp, Zhong:2024qqs, Cheng:2024lsn] and experimental measurements [CLEO:1995qyd], and shadows representing the distribution of $|A|$ and $|B|$ obtained in this work, $G_{F}$ is Fermi coupling constant, $c$ and $d$ means whether the model uses the results from Belle experiment with the BFs and decay asymmetry parameters of $\Xi_{c}^{0}\to\Xi^{0}\pi^{0}$, $\Xi^{0}\eta$, and $\Xi^{0}\eta^{\prime}$. The contours correspond to 68.2%, 95.4%, 99.7% confidence levels, respectively.
@@
A secondary fit is also performed where the CPC\!P constraint is removed and is used to test the SM, and the results are shown in Appendix B. The $\alpha$-induced CPC\!P observable, weak and strong phase shifts are constructed in $A^{\alpha_{BP}}_{C\!P}=\frac{\alpha_{BP}+\bar{\alpha}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, $\text{tan}\phi_{C\!P}=\frac{\beta_{BP}+\bar{\beta}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, $\text{tan}\Delta_{S}=\frac{\beta_{BP}-\bar{\beta}_{BP}}{\alpha_{BP}-\bar{\alpha}_{BP}}$, where $\beta_{BP}=\sqrt{1-\alpha_{BP}^{2}}\sin\Delta_{BP}$ and $\bar{\beta}_{BP}=\sqrt{1-\bar{\alpha}_{BP}^{2}}\sin\bar{\Delta}_{BP}$. The first two observables, $A^{\alpha_{BP}}_{C\!P}$ and $\tan\phi_{C\!P}$, are CPC\!P-violating quantities, corresponding to the two constraints $\alpha_{BP}=-\bar{\alpha}_{BP}$ and $\Delta_{BP}=-\bar{\Delta}_{BP}$ imposed in the CPC\!P-conserving baseline fit, while $\tan\Delta_{S}$ is a CPC\!P-conserving observable. The $\phi_{C\!P}$ and $\Delta_{S}$ denote the weak interaction phase shift between $P$-wave and $S$-wave introduced by phase angle in the Cabibbo-Kobayashi-Maskawa (CKM) matrix and the strong phase shift. The results are summarized in Table 4. Some of the parameters in the table are derived from the fit results, and their uncertainties account for correlations among the fit parameters. The full correlation matrix is provided in Table A2 of Appendix A.
@@
Table 4: Results of decay observables, where $\langle\alpha_{BP}\rangle$, $\langle\Delta_{BP}\rangle$, and $\delta_{p}-\delta_{s}$ are measured based on the assumption of CPC\!P conservation, while others are under the assumption of CPC\!P violation. Terms that could not be extracted are denoted by “”. All angular values in the table are expressed in radians. Among the eight parameters, $\langle\alpha_{BP}\rangle$ and $\langle\Delta_{BP}\rangle$ are directly obtained from the fit, $\langle\beta_{BP}\rangle$, $\langle\gamma_{BP}\rangle$, and $\delta_{p}-\delta_{s}$ are quantities derived from them. And $A^{\alpha_{BP}}_{C\!P}$, $\text{tan}\phi_{C\!P}$, and $\text{tan}\Delta_{S}$ are derived from the results given in Appendix B.
Parameter $\phantom{-}\Lambda_{c}^{+}\to pK_{S}^{0}$ $\phantom{-}\Lambda_{c}^{+}\to\Lambda\pi^{+}$ $\phantom{-}\Lambda_{c}^{+}\to\Sigma^{0}\pi^{+}$ $\phantom{-}\Lambda_{c}^{+}\to\Sigma^{+}\pi^{0}$